LearnCo manufactures and sells one product, an abacus for classroom use, with two models, the Basic model and the Deluxe model. The company began operations on January 1, 20Y1, and is planning for 20Y2, its second year of operations, by preparing budgets from its master budget.

The company is trying to decide how many units to manufacture, how much it might spend on direct materials and direct labor, and what their factory overhead expenses might be. In addition, the company is interested in budgeting for selling and administrative costs, and in creating a budgeted income statement showing a prediction of net income for 20Y2.

You have been asked to assist the controller of LearnCo in preparing the 20Y2 budgets.

The sales budget often uses the prior year’s sales as a starting point, and then sales quantities are revised for various factors such as planned advertising and promotion, projected pricing changes, and expected industry and general economic conditions. LearnCo has completed reviewing its prior year’s sales and has prepared the following sales budget.

After reviewing LearnCo’s sales budget, you note that three numbers have been omitted. The company’s controller has told you that the units sold for the Basic and Deluxe models are expected to be the same. Fill in the missing amounts.

LearnCo
Sales Budget
For the Year Ending December 31, 20Y2
Product Unit Sales Volume Unit Selling Price Total Sales
Basic Abacus $8.00 $288,000
Deluxe Abacus 432,000
Totals 72,000 $720,000
Answer #1

--Requirement, as asked

LearnCo
Sales Budget
For the Year Ending December 31, 20Y2
Product Unit Sales Volume Unit Selling Price Total Sales
Basic Abacus                36,000 $8 $288,000
Deluxe Abacus                36,000 $12 $432,000
Totals                72,000 $720,000

--Calculations

LearnCo
Sales Budget
For the Year Ending December 31, 20Y2
Product Unit Sales Volume Unit Selling Price Total Sales
Basic Abacus =$288000/$8 8 288000
Deluxe Abacus =72000-36000 =$432000/36000 432000
Totals 72000 =288000+432000